Displaying report 1-1 of 1.
Reports until 11:57, Tuesday 22 April 2025
GQuEST Output Filter Cavities
Torrey Cullen - posted 11:57, Tuesday 22 April 2025 - last comment - 12:15, Tuesday 22 April 2025(12265)
OFC3 Alignment complete

OFC3 1550 alignment is complete. To recap, the 1550 OFC3 path has no test light. It is aligned directly to the output of OFC2. In order to achieve this we built a 1550 PDH scheme for OFC2 for more stable locking. I can then scan on the piezo in OFC3 to find cavity flashes. Additionally I had to implement some periscoping mirrors as outlined in previous posts.

One problem: OFC3.png is a cavity scan. In red is the PDH error signal for the 775 light and in blue is the 1550 TRANS signal. Fairly clean cavity scan. There seems to be a large unexplained frequency drift, as seen in ScreenRecording2025-04-22114505.mp4. This drift is significant portion of the cavity FSR (10's of MHz) and appears to just be the cavity/piezo. I don't believe it is the laser, as you can track the correction voltage being applied to the DC Modulation port of the teraxion via its interface teraxion.png. May be worth recording this data and making sure but quick look by eye it doesn't seem to be the case. Additional proof for this not being the laser is the fact the 775 and 1550 drift the same amount. If it was the laser the 775 light would be moving twice as fast due to the SHG frequency doubling.

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 12:15, Tuesday 22 April 2025 (12266)

After talking with Sander I think these may make sense. These are bigger than the normal fluctuations seen in the other cavities, but there are still some in the other cavities. The explanation being, the first cavity has some mechanical fluctuations that the laser has to track. Since we lock the first cavity to the laser, those fluctuations are imprinted on the laser light that is exiting the first cavity. Then the new cavity has its own fluctuations PLUS the first cavity fluctuations to track with the piezo. I think this explains the larger fluctuations in the new cavity.

Displaying report 1-1 of 1.